OpenStudy (worldwarkatyy):

What is a simpler form of the radical expression? 4 to the square root of 625x^8y^12 A. 5x^2|y^3| B.25|x^2|y^3| C. 5x^4|y^3| D. 25|x^4y^3

OpenStudy (worldwarkatyy):

Help plsssss

OpenStudy (worldwarkatyy):

WILL GIVE A MEDAL TO SOMEONE WHO CAN ANSWER

jhonyy9 (jhonyy9):

625=25^2 x^8 =(x^4)^2 y^12 =(y^6)^2 <<4 to the squareroot >> what mean exactly this 4 ?